OpenStudy (gabylovesyou):

Choose the equation below that represents the line passing through the point (-5, 1) with a slope of 3/2. y - 5 = 3/2(x + 1) y + 1 = 3/2(x - 5) y + 5 = -3/2(x - 1) y - 1 = 3/2(x + 5)

OpenStudy (compassionate):

Didn't we just go over this? c:

OpenStudy (gabylovesyou):

just wanted to make sure my answer was correct lol

OpenStudy (compassionate):

First you need to put these equations in y=mx+b form.

OpenStudy (anonymous):

\[y-y1=m(x-x1)\]

OpenStudy (compassionate):

\[y = \frac{ 3x + 13 }{ 2 }\] \[y=\frac{ 3x + 17 }{ 2 }\] \[y=\frac{ -3x -7 }{ 2 }\]

OpenStudy (compassionate):

No, @monokerous , we're doing things my way D:<

OpenStudy (ghazi):

y-1=3/2(x+5)

OpenStudy (anonymous):

where m=3/2, y1=1 and x1=-5 Just an alternative.

OpenStudy (compassionate):

You guys ruined everything DX

OpenStudy (anonymous):

@Compassionate , sorry... I just wanted to point out that there was an easier way. D: Didn't mean to ruin anything!

OpenStudy (gabylovesyou):

lol is it D??

OpenStudy (compassionate):

:'(

OpenStudy (anonymous):

Feast on a medal and be merry? :)

OpenStudy (anonymous):

Yes it is.

OpenStudy (compassionate):

*nomnomnom.*

OpenStudy (gabylovesyou):

k thanks

OpenStudy (anonymous):

you're welcome.